What is the equilibrium Zn2+ concentration when 1.78 L of a 0.160 M zinc fluoride solution are mixed with 1.94 L of a 0.183 M ammonium phosphate solution? [Zn2+] = _______M
Answer #1

1) Ksp of Zn3(PO4)2 = 9*10^-33

Zn3(PO4)2(s) ----> 3Zn+2(aq) + 2PO4-3(aq)

Ksp of Zn3(PO4)2 = (3s)^3*(2s)^2

9*10^-33 = (3s)^3*(2s)^2

S = solubility of Zn3(PO4)2 = 1.53*10^-7 mole/l

Concentration of Zn+3 ion = 3s

                          = 3*1.53*10^-7

                          = 4.59*10^-7 M